OpenStudy (lorenzo):

Use 2,3,5,8 once with combo of +, -, mutiplication, division, or ( ) with value if 104. Can repeat operations but not numbers. answer?

OpenStudy (anonymous):

One way to do these is look for factors. 104 = 8*13 so guess that this is where 8 is used. We then need to make 13 using 2,3,5. 13 is prime so addition must have been used to form it, from this we expect a multiple and addition. 13 = 2*5 + 3 so: 104 = 8*(3+2*5)